A sinusoidal voltage v(t) = (40 V) sin(100t) is ap
Answer #1

Xl = Lw = 160*10^-3 * 100 =   16 ohm

Xc = 1/cw = 1/( 99*10^-6*100) =101.010101 ohm

part a)

Z = sqrt[ (101.0101-16)^2+68^2] =108.860999 ohm

Part b)

current Im = V/z =40 / 108.860999 =0.367441052

Part c)

Im = 40 / 108.860999 =0.367441052

w = 100

phi =atan ( ( 101-16) /68) = 51.34 degree